OpenStudy (icalibear):

Which equation is equivalent to x^2 + x = 12? a) (x - 4)(x+3)=0 b) (x+4)(x-3)=0 c) (x+12)(x+1)=0 d) (x-12)(x-1) = 0

OpenStudy (anonymous):

B, would you like me to explain how i got it?

OpenStudy (icalibear):

Yes, that was my first answer and I broke it down into x^2 - 12 = 0 x^2 = (x )(x ) = 0 4(-3) = -12 (x + 4)(x - 3) = 0 But that may be incorrect so please explain

OpenStudy (anonymous):

What i do first is that I look at the middle number \[x ^{2}+ x -12= 0\], you have to make the equation equal 0. the middle number is X. 1. you would look at your answers. 2. you add the (x+ 4)(X-3) 4+3= 1 3.than you 4 X -3 = -12 that is when the equation is what i did on top

OpenStudy (anonymous):

and for 2. it would be 4x+3x= X

OpenStudy (icalibear):

Okay. But why couldnt it just as easily be A, if -3(4) and (-4)3 are both -12?

OpenStudy (anonymous):

-4x +3 =